Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2014-04-30

Observers: Sandra Benitez, Joe Polshaw
Support: Kate Maguire 
TAT on duty: Andrea Pastorello, Cosimo Inserra

--------------------------------------------------------------

WEATHER CONDITIONS:
start: wind 5 m/s; temp. 8.3 C; humid. 35%; Some clouds on the horizon, almost entirely clear sky 
middle: wind 8.3 m/s; temp. 9.9 C: humid. 14%; Clear, became cloudy at about 6:00 UT
end: wind 10.4 m/s; temp. 10.5; humid. 10%; Cloudy


--------------------------------------------------------------

INSTRUMENTS: EFOSC2, SOFI

CALIBRATIONS

Afternoon:
-- EFOSC bias, arcs, dome flats

Twilight:
-- Sky flats: UB -- Too late - will repeat in the morning. (or not - it was cloudy!)


--------------------------------------------------------------

OBSERVATIONS

Target                   Type                         UT-Start         Filter/Grism           Slit (")         Seeing (")      Notes

*** EFOSC2 ***

OGLE-016             Follow (IMAG)            23:07             VRi                        N/A             1                  (1)
l745a                    STD (SPEC)                 00:40             Gr11/13/16         1/10             0.9              (2)
CSS140424...        Follow (SPEC)             01:10             Gr11/16                1                  1                 (3)
CSS140424...        Follow (IMAG)            02:59            UBVRi                    N/A               1                 (4)
CSS140421...        Follow (SPEC)            03:25             Gr13                      1                   0.9             (5)
CSS140421...        Follow (IMAG)            05:32             UBVRi                    N/A               1                 (6)

*** Change to SOFI ***
*** Cloud cover across the whole sky during the switch to SOFI until the end of the night ***



(1) OGLE-2013-SN-016. Templates in VRi, 6*100s each. Used "EFOSC_img_acq_PresetNoGuide" to acquire the target due to previous problems with finding a guide star. First two exposures started too early (sky too bright). Since the target was low and declining, we did not perform image analysis but just did a focus.
(2) Full image analysis. We lost quite a lot of time here since there was a problem with the telescope, which took a while to identify and fix. Clear sky so we used slit 10" instead of 1.5" in order to test the throughput. We will do 1"/1.5"/10" with the STD later in the night (if all goes well), since there are a lot of targets to observe early in the night.
(3) CSS140424-133007-212728. 2*1800s gr11, 1800s gr16. Very interesting and decent S/N spectra. H-alpha in emission only (z~0.018). In the blue, it looks as though there are many low velocity (~1500 km/s ?) p-cygni lines, including more of the Balmer series.
(4) CSS140424-133007-212728. U 300s, B 45s, V 30s, Ri 20s
(5) CSS140421-142042+031602. Gr13, 3*2400s
(6) CSS140421-142042+031602. U 3*300s, B 200s, V 90s, Ri 60s

Comments